OpenStudy (anonymous):

To win a lottery, a person must pick five numbers from 1 to 45 in any order AND a MEGA number from 1 to 45 on the play slip. Determine the probability for winning the jackpot. A. 0 B. 1/54,979,155 C. 1/146,611,080 D. 1/6,597,498,600

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Jim766

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (poopsiedoodle):

I have no idea o.o

OpenStudy (anonymous):

I know its Diffenitly not D

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

I'm horrible at probability sorry :( the only way I would have done it was (options)^(slots) 45^6 = 8303765625 but that's not an option, so I don't know, sorry :(

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (jim766):

we have to multiply all the options together for each number how many options are there when they go to select the first number?

OpenStudy (anonymous):

5 options

OpenStudy (jim766):

the number choice for the first slot is 1 through 45, how many numbers is that?

jigglypuff314 (jigglypuff314):

oh the numbers can't be reused? :/ that'd help (total options) * (total minus 1) * (total minus 2) etc for number of options

OpenStudy (jim766):

yes once a number is selected, it cannot be selected again

OpenStudy (anonymous):

1-45.. 45 Numbers? .... Shouldnt it be like 45!/1!/44! ?

OpenStudy (jim766):

1st number _______ 45 possible numbers 2nd number _______ 44 possible numbers

OpenStudy (jim766):

3rd number is 43 possible numbers 4th number is 42 possible numbers 5th numbrer is 41 possible numbers

OpenStudy (jim766):

so...for the first 5 numbers 45*44*43*42*41 but the 6th number is a mega number, now we can choose from 1-45 again

jigglypuff314 (jigglypuff314):

yes, but I don't understand why is not D if you do it that way...

OpenStudy (anonymous):

146,611,080

OpenStudy (jim766):

so it would be 45*44*43*42*41*46 that will give you the total number of possible lottery #s

OpenStudy (anonymous):

I picked D the first time and it was wrong @jigglypuff314

OpenStudy (anonymous):

6,744,109,680 @Jim766

OpenStudy (jim766):

I would say it is D, you sure that it is wrong?

jigglypuff314 (jigglypuff314):

this is a weird lottery... they win if they manage to pick numbers? it didn't say what you needed to win... are you missing part of the question? if you need the mega to win 45*44*43*42*41*45 = 6597498600 if you don't need mega to win 45*44*43*42*41 = 146611080

OpenStudy (jim766):

1 winner ------------------ 6,597,498,600 possible #s

OpenStudy (anonymous):

yes im sure my teacher said it was wrong also

jigglypuff314 (jigglypuff314):

then you probably don't need the mega to win...? the question is phrased oddly, if never really says how someone wins just says they win by... picking numbers... are you sure that's the full question?

OpenStudy (anonymous):

yes im sure

jigglypuff314 (jigglypuff314):

if you don't need mega to win 45*44*43*42*41 = 146611080 that's all I've got...

OpenStudy (anonymous):

thanks for yall help , I appreciate it.

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

I don't like probability or stats i choose c 1/45*1/44*1/43*1/42*1/41= 1/146611080

OpenStudy (anonymous):

thank you @triciaal

OpenStudy (triciaal):

i'm not sure i am thinking that the MEGA number is already accounted for on the slip 5/45

OpenStudy (jim766):

no I think there are 6 numbers all together , 6th is the mega

OpenStudy (jim766):

the reason I think so is that it very clearly states "and"

OpenStudy (triciaal):

i think the number is already on the slip so it would be one of the five numbers already chosen I am acknowledging the word "and"